OpenStudy (anonymous):

(2/3)^2 + 1/7 I know we solve for the parenthesis first then the ^2 and finally add .. but 2/3 = ?

OpenStudy (anonymous):

this question is to add the two yes?

OpenStudy (anonymous):

\[(\frac{a}{b})^2=\frac{a}{b}\times \frac{a}{b}=\frac{a\times a}{b\times b}=\frac{a^2}{b^2}\] so it is straightforward to square a fraction, square top and bottom

OpenStudy (anonymous):

sry for disturbing...

OpenStudy (anonymous):

satellite 37 pls help me after solve this problem.. thx...

OpenStudy (anonymous):

thus \[\left(\frac{2}{3}\right)^2=\frac{2^2}{3^2}=\frac{4}{9}\]

OpenStudy (anonymous):

so the total is 5/16 with the 1/7 included?

OpenStudy (anonymous):

\[\frac{4}{9}+\frac{1}{7}=\frac{7\times 4+9\times 1}{7\times 9}\]

OpenStudy (anonymous):

you do not add top and bottom separately when you add fractions

OpenStudy (anonymous):

so 37/63?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much